# Empowering Robot Path Planning with Large Language Models: osmAG Map Topology & Hierarchy Comprehension with LLMs

### v2

@@ I Introduction
Recent years have seen a growing interest in Large Language Models (LLMs) like ChatGPT[1] and LLaMA[2]. Real-life robots often face unpredictable situations, such as a campus delivery robot blocked by a closed intersection for pipe repair, depicted in Fig. 1. Despite the construction notice being posted publicly, the robot was unaware. Integrating general knowledge with real-time data from public notice boards through LLMs could significantly improve navigation and decision-making. For this integration to be effective, the robot’s ’brain’ needs to grasp the map’s hierarchy and topology.
@@ II-A Integrated Robotics and Navigation with LLMs and Maps
Research in robotics integrates natural language models like PaLM-E[10], which uses real-world sensor data for better decision-making, and LLMs for tasks such as code generation[11], object rearrangement[12], and motion planning[13]. Visual-language navigation systems like [14] allow robots to follow human instructions using advanced map technologies such as VLMaps[15] that combine language models with 3D point clouds. Our focus is on developing map representations that are easily understandable by both LLMs and robots, prioritizing map data to enhance navigation and decision-making in robotics.
@@ III-A Prompt Engineering
As shown in Fig. 4 and referenced in [3], the original Area Graph uses ’passage’ to connect different areas, with ’passage’ and ’area’ under separate ’way’ tags in the osmAG XML, complicating connection understanding for LLMs. We developed a script to create osmAG variants that integrate connection details directly within the ’area’ tags. Variant 1 adds a ’connected_area’ key directly to the ’area’, specifying its connected area. Variant 2 further specifies the current area’s name and its direct connections.
Figure 4: The osmAG map representation and two of its variants. The osmAG enclosed in blue box represents the original format that uses ’passage’ to describe connections between areas. In the yellow box we illustrate Variant 1 of osmAG, which introduces a tag with a key set to ”connected_area” and a value corresponding to the area connected via the passage. Variant 2 of osmAG displayed in the green box, modifies Variant 1 by replacing ”connected_area” with ”current area name_directly_connected_room”.
@@ III-C1 Topological Datasets
Map templates shown in Fig. 5 are handcrafted using JOSM, with ’area’ and ’passage’ defined per osmAG standards in [6]. Labels like ’1d-201’ indicate a room’s location by zone, floor, wing, and room number. We employ a script to randomize these attributes to create varied maps from a single template. Room numbers are shuffled to ensure paths rely on map information, not sequence. For each map, we ask the LLM to find a path (sequence of room names) between each two rooms as one item of the dataset. Dataset 1, using the ’normal’ layout Template (a) from Fig. 5, contains 440 entries. Dataset 2, mixing Templates (a), (b), and (c), includes 12,520 entries aimed for training, with 440 reserved for testing. Dataset 3, derived from an Template (d) that is not used in the training dataset, aim to evaluate the model’s generalization ability. The number of rooms in templates is limited by token limits, preventing larger designs. However, leaf rooms with a single door are unnecessary for path planning unless they are the terminal rooms, so they can be omitted beforehand. Thus, despite the template’s limited number of rooms, it remains highly relevant for real-world, large environment applications.
@@ III-C2 Hierarchical Datasets
As shown in Fig. 6, we randomly selected two osmAG from Dataset 2, assigning each room a random ’owner’ tag. We structured each map with ’parent’ tags linking rooms to zones, floors, and wings, and assigned buildings ’SIST_1’ or ’SIST_2’ to establish a hierarchy. This setup supports queries asking the LLM to locate individuals by directing towards the correct building. We created 1056 such queries for training Dataset 4. An example of this map visualized by JOSM is in Fig. 6. The LLM must navigate hierarchy tree to correctly identify the target building.
@@ III-C3 General Knowledge Dataset
LoRA fine-tuning is not completely immune to catastrophic forgetting [21]. To verify that general knowledge capabilities remain intact, we created a small Dataset 5 with 20 general questions, like ”Who wrote ’Hamlet’?”, to test the model after fine-tuning.
Figure 6: This image displays two osmAG maps from Dataset 2, each randomly assigned a ’owner’ tag per room (first names shown for clarity). The maps are structured by zone, floor, and wing, with buildings ’SIST_1’ or ’SIST_2’ to establish a hierarchical tree using ’parent’ tags. Details for the left map are: SIST_2, 3_wing, 7_floor, b_zone.
@@ III-D Fine-tuning
We utilize Meta’s LLaMA2 model[2], a decoder-only transformer, as our base model. The models are fine-tuned using datasets from Section III-C with the LoRA (Low Rank Adaptation)[19] method, which updates additive low-rank matrices in neural layers while keeping the original weights frozen, reducing the number of trainable parameters and allowing efficient use of computational resources. Fine-tuning targets topological tasks with Dataset 2 and hierarchical tasks with Dataset 4. The LoRA hyperparameters are as follows: the rank is set to 8, and target modules of LoRA are set to ”q_proj” and ”v_proj”. These two modules are the query and value matrices in the self-attention mechanism of the transformer architecture [19, 25]. The learning rate is set to 5e-5 and we opted for a cosine learning rate scheduler. The fine-tuning process is illustrated in Figure 7. The LLaMA2-7B and LLaMA2-13B models were fine-tuned for topological and hierarchical tasks using 4$\times$×NVIDIA A40 and 4$\times$×A100 GPUs, respectively. Topological tasks took 6.6 hours for LLaMA2-7B and 5.5 hours for LLaMA2-13B, both running 2.5 epochs with a dataset of 12,520 entries. Hierarchical tasks required 2 hours for LLaMA2-7B and 1.5 hours for LLaMA2-13B, each undergoing 7.76 epochs with 1,056 entries.
@@ IV Experiments
IV-A Prompt Engineering Experiment
IV-B osmAG Variant Experiment

### v3

@@ I Introduction
Recent years have seen a growing interest in Large Language Models (LLMs) like ChatGPT[1] and LLaMA[2]. Real-life robots often face unpredictable situations, such as a campus delivery robot blocked by a closed intersection for pipe repair, depicted in Fig. 1. Despite the construction notice being posted publicly, the robot was unaware. Integrating general knowledge with real-time data from public notice boards through LLMs could significantly improve navigation and decision-making. For this integration to be effective, the robot’s ‘brain’ needs to grasp the map’s hierarchy and topology.
@@ II-A Integrated Robotics and Navigation with LLMs and Maps
Research in robotics integrates natural language models like PaLM-E[10], which uses real-world sensor data for better decision-making, and LLMs for tasks such as code generation[11], object rearrangement[12], and motion planning[13]. Visual-language navigation systems like [14] allow robots to follow human instructions using advanced map technologies such as VLMaps[15] that combine language models with 3D point clouds. [16] integrates external information into robot navigation by utilizing osmAG as map representation. Our focus is on developing map representations that are easily understandable by both LLMs and robots, prioritizing map data to enhance navigation and decision-making in robotics.
@@ III-A Prompt Engineering
As shown in Fig. 4 and referenced in [4], the original Area Graph uses ‘passage’ to connect different areas, with ‘passage’ and ‘area’ under separate ‘way’ tags in the osmAG XML, complicating connection understanding for LLMs. We developed a script to create osmAG variants that integrate connection details directly within the ‘area’ tags. Variant 1 adds a ‘connected_area’ key directly to the ‘area’, specifying its connected area. Variant 2 further specifies the current area’s name and its direct connections.
Figure 4: The osmAG map representation and two of its variants. The osmAG enclosed in blue box represents the original format that uses ‘passage’ to describe connections between areas. In the yellow box we illustrate Variant 1 of osmAG, which introduces a tag with a key set to “connected_area” and a value corresponding to the area connected via the passage. Variant 2 of osmAG displayed in the green box, modifies Variant 1 by replacing “connected_area” with “current area name_directly_connected_room”.
@@ III-C1 Topological Datasets
Map templates shown in Fig. 5 are handcrafted using JOSM, with ‘area’ and ‘passage’ defined per osmAG standards in [3]. Labels like ‘1d-201’ indicate a room’s location by zone, floor, wing, and room number. We employ a script to randomize these attributes to create varied maps from a single template. Room numbers are shuffled to ensure paths rely on map information, not sequence. For each map, we ask the LLM to find a path (sequence of room names) between each two rooms as one item of the dataset. Dataset 1, using the ‘normal’ layout Template (a) from Fig. 5, contains 440 entries. Dataset 2, mixing Templates (a), (b), and (c), includes 12,520 entries aimed for training, with 440 reserved for testing. Dataset 3, derived from an Template (d) that is not used in the training dataset, aim to evaluate the model’s generalization ability. The number of rooms in templates is limited by token limits, preventing larger designs. However, leaf rooms with a single door are unnecessary for path planning unless they are the terminal rooms, so they can be omitted beforehand. Thus, despite the template’s limited number of rooms, it remains highly relevant for real-world, large environment applications.
@@ III-C2 Hierarchical Datasets
As shown in Fig. 6, we randomly selected two osmAG from Dataset 2, assigning each room a random ‘owner’ tag. We structured each map with ‘parent’ tags linking rooms to zones, floors, and wings, and assigned buildings ‘SIST_1’ or ‘SIST_2’ to establish a hierarchy. This setup supports queries asking the LLM to locate individuals by directing towards the correct building. We created 1056 such queries for training Dataset 4. An example of this map visualized by JOSM is in Fig. 6. The LLM must navigate hierarchy tree to correctly identify the target building.
@@ III-C3 General Knowledge Dataset
LoRA fine-tuning is not completely immune to catastrophic forgetting [22]. To verify that general knowledge capabilities remain intact, we created a small Dataset 5 with 20 general questions, like “Who wrote ‘Hamlet’?”, to test the model after fine-tuning.
Figure 6: This image displays two osmAG maps from Dataset 2, each randomly assigned a ‘owner’ tag per room (first names shown for clarity). The maps are structured by zone, floor, and wing, with buildings ‘SIST_1’ or ‘SIST_2’ to establish a hierarchical tree using ‘parent’ tags. Details for the left map are: SIST_2, 3_wing, 7_floor, b_zone.
@@ III-D Fine-tuning
We utilize Meta’s LLaMA2 model[2], a decoder-only transformer, as our base model. The models are fine-tuned using datasets from Section III-C with the LoRA (Low Rank Adaptation)[20] method, which updates additive low-rank matrices in neural layers while keeping the original weights frozen, reducing the number of trainable parameters and allowing efficient use of computational resources. Fine-tuning targets topological tasks with Dataset 2 and hierarchical tasks with Dataset 4. The LoRA hyperparameters are as follows: the rank is set to 8, and target modules of LoRA are set to “q_proj” and “v_proj”. These two modules are the query and value matrices in the self-attention mechanism of the transformer architecture [20, 23]. The learning rate is set to 5e-5 and we opted for a cosine learning rate scheduler. The fine-tuning process is illustrated in Figure 7. The LLaMA2-7B and LLaMA2-13B models were fine-tuned for topological and hierarchical tasks using 4$\times$×NVIDIA A40 and 4$\times$×A100 GPUs, respectively. Topological tasks took 6.6 hours for LLaMA2-7B and 5.5 hours for LLaMA2-13B, both running 2.5 epochs with a dataset of 12,520 entries. Hierarchical tasks required 2 hours for LLaMA2-7B and 1.5 hours for LLaMA2-13B, each undergoing 7.76 epochs with 1,056 entries.
@@ IV Experiments
IV-A Prompt Engineering Experiment
IV-B osmAG Variant Experiment